OpenStudy (anonymous):

What is the relationship between the lines determined by these two equations? 3x + 6y = 8 y = 2x – 8

OpenStudy (anonymous):

try to get the lines in the same form. so for the first one, 3x+6y = 8 solve for y 6y = -3x + 8 y = -(3/6)*x + 8/6 y = -(1/2)x = 4/3 Notice now that the slope of this line is -1/2. The slope of the second line (y = 2x - 8) is 2 We know this because y= mx +b; m= slope When the slopes are the negative reciprocals of one another, that means that they are perpendicular!

OpenStudy (anonymous):

thank you